OpenStudy (anonymous):

Substitute y= 2x+4 3x+4y=5

OpenStudy (anonymous):

y= 2x+4 3x+4y=5

OpenStudy (anonymous):

just replace y with 2x+4 everywhere you see it 3x + 4(2x+4) = 5 what is 4(2x+4)?

OpenStudy (anonymous):

want me to substitute 4(2x+4)?

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

distribute the 4 and expand it

OpenStudy (anonymous):

9x+16

OpenStudy (anonymous):

8* not 9

OpenStudy (anonymous):

i need to get X and Y

OpenStudy (anonymous):

8x+16* yeah So now we have 3x + 8x+16 = 5 Combine like terms. 11x + 16 = 5 now what?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

yeah and to find y replace -1 into y= 2x+4

OpenStudy (anonymous):

so y=2(-1)x+4?

OpenStudy (anonymous):

basically and 2*(-1)+4 = -2 + 4 = 2

OpenStudy (anonymous):

y=4−2x

OpenStudy (anonymous):

the whole point is x = -1. You basically have y. y=2(-1)+4? NO X'S HERE y = -2 + 4 = 2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@soisoi

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

possibly? post it in another question